OpenStudy (anonymous):

Solve for x: -3(x + 3) = -3(x + 1) - 5. (1 point) 6 -6 All real numbers No solution

OpenStudy (smored):

\[-3(x+3)=-3(x+1)-5\]distribute both sides\[-3x-9=-3x-3-5\]\[-3x-9=-3x-8\]\[-3x=-3x+1\]\[0=1\]there is no solution.

OpenStudy (anonymous):

thank you hun :)